For nominal interest rate of compounding is continuous. Show your solution 3. %, effective annual interest rate will be 12% when 4. Under what conditions APR will be different from EAR? L. Shark is designing a new account that pays interest quarterly. They wish to pay effectively, a 16% per year on this account. L. Shark desires to advertise the annual percentage rate on this new account (and not the effective rate, since their competitors state their interest on an annualized basis). what is the APR that corresponds to an effective rate of 16% for this new account? 5.
Answer #1

3:

APR = In( EAR + 1)

= In (1.12)

= 11.33%

4: APR will be different from EAR when there are multiple compounding in a year.

5: EAR = 16%

N = 4

EAR = (1+ APR/m)^m -1

0.16 = (1+ APR/4)^4 -1

APR = 15.12%